# Knowledge-to-SQL: Enhancing SQL Generation with Data Expert LLM

### v2

@@ 7 Ethics Statement
We discuss the implementation details of our proposed framework. We used parameter-efficient fine-tuning (PEFT) in the training stage to train our models. Specifically, in each stage (supervised fine-tuning and DPO preference learning), we utilize low-rank adaptation (LoRA) Hu et al. (2021) as PEFT method, the trainable parameters occupy 0.0622% of full parameters. All the experiments are conducted on four NVIDIA A800-SXM4-80GB GPUs, and the transformers package version is 4.36.2. The details of training and hyper-parameters are as follows.
@@ A.1 Supervised Fine-tuning
As previous introduced, As previously introduced, the base model of DELLM we selected is LLaMA-2-13b with official configuration [^1] https://huggingface.co/meta-llama; the training details are listed in Table 5. The model’s architecture is identical to the official provided in Huggingface.
@@
Table 5: Hyper-parameters of the SFT stage of DELLM’s training
@@ A.2 DPO Training
The hyper-parameters of DPO is similar with the previous stage, the details are shown in Table 6.
@@
Table 6: Hyper-parameters of the DPO training for DELLM.
@@ A.3 Generation Configuration
In each generation parts of our framework during training, the configuration are identical. The details are listed in Table 7.
@@ A.4 Versions of Proprietary LLMs
The version of ChatGPT in this work is GPT-3.5-Turbo-1106  OpenAI (2023), and version of GPT-4 is GPT-4-0613  OpenAI (2023).

### v3

@@ Acknowledgements
We discuss the implementation details of our proposed framework. We used parameter-efficient fine-tuning (PEFT) in the training stage to train our models. Specifically, in each stage (supervised fine-tuning and DPO preference learning), we utilize low-rank adaptation (LoRA) Hu et al. (2021) as PEFT method, the trainable parameters occupy 0.0622% of full parameters. All the experiments are conducted on four NVIDIA A800-SXM4-80GB GPUs, and the transformers package version is 4.36.2. The details of training and hyper-parameters are listed as follows.
@@ A.1 Supervised Fine-tuning
As previously introduced, the backbone model of DELLM is selected as LLaMA-2-13b with official configuration [^4] https://huggingface.co/meta-llama; the training details are given in Table 6. The model’s architecture is identical to the official provided in Huggingface.
@@
Table 6: Hyper-parameters of SFT.
@@ A.2 DPO Training
The hyper-parameters of DPO are similar to the previous stage, the details are shown in Table 7.
@@
Table 7: Hyper-parameters of the DPO.
@@ A.3 Generation Configuration
In each generation of parts of our framework during training and testing, the configuration is identical. The details are listed in Table 8.
@@ A.4 Versions of Proprietary LLMs
The version of ChatGPT in this work is GPT-3.5-Turbo-1106, and the version of GPT-4 is GPT-4-0613.